# Federated Domain Generalization with Data-free On-server Gradient Matching

### v1

@@ Abstract
Domain Generalization (DG) aims to learn from multiple known source domains a model that can generalize well to unknown target domains. One of the key approaches in DG is training an encoder which generates domain-invariant representations. However, this approach is not applicable in Federated Domain Generalization (FDG), where data from various domains are distributed across different clients. In this paper, we introduce a novel approach, dubbed Federated Learning via On-server Matching Gradient (FedOMG), which can efficiently leverage domain information from distributed domains. Specifically, we utilize the local gradients as information about the distributed models to find an invariant gradient direction across all domains through gradient inner product maximization. The advantages are two-fold: 1) FedOMG can aggregate the characteristics of distributed models on the centralized server without incurring any additional communication cost, and 2) FedOMG is orthogonal to many existing FL/FDG methods, allowing for additional performance improvements by being seamlessly integrated with them. Extensive experimental evaluations on various settings to demonstrate the robustness of FedOMG compared to other FL/FDG baselines. Our method outperforms recent SOTA baselines on four FL benchmark datasets (MNIST, EMNIST, CIFAR-10, and CIFAR-100), and three FDG benchmark datasets (PACS, VLCS, and OfficeHome).
@@ 1 Introduction
Federated Learning (FL) has gained widespread recognition due to its ability to train models collaboratively across multiple clients while keeping their individual data secure. However, one of the practical problems, how to make models trained on sites of heterogeneous distributions generalize to target clients of unknown distributions, i.e., Federated Domain Generalization (FDG), remains under explored. While label distribution shift has been considered in traditional FL, FDG focuses on the feature shift among clients and considers each client as an individual domain.
FDG shares a similar goal as standard Domain Generalization (DG) (Nguyen et al., 2022b; Li et al., 2018), i.e., generalizing from multi-source domains to unseen domains. However, unlike DG, where knowledge from various domains can be jointly utilized to develop an efficient algorithm, FDG prohibits direct data sharing among clients, which makes most existing DG methods hardly applicable. To address the challenges inherent to FDG, recent methods circumvent these difficulties by adopting alternative strategies, e.g., unbiased local training within each isolated domain, local data preprocessing Huang et al. (2022; 2024), or knowledge sharing among users (Liu et al., 2021; Chen et al., 2023). As a consequence, an open question remains for FDG:
@@
    Due to the design of on-server gradient matching, our method is orthogonal to many classic FDG methods, where current FDG methods focus on the local training Huang et al. (2022; 2024). As a consequence, the integration of FedOMG with other FDG methods can achieve improved performance.
@@ 2 Preliminaries
Problem Setup: We consider an FL system comprising a set of source clients ${\mathcal{U}}_{\mathcal{S}}=\{u|u=1,2,\ldots,U_{\mathcal{S}}\}$, and target clients ${\mathcal{U}}_{\mathcal{T}}=\{u|u=U_{\mathcal{S}}+1,2,\ldots,U_{\mathcal{S}}+U_{\mathcal{T}}\}$. Each client $u$ gains access to its data ${\mathcal{D}}_{u}=\{x_{i},y_{i}\}^{N_{u}}_{i=1}$ with $N_{u}$ data samples, where $u\in\{{\mathcal{U}}_{\mathcal{S}};{\mathcal{U}}_{\mathcal{T}}\}$. Source and target clients are assigned to the source datasets ${\mathcal{D}}_{\mathcal{S}}=\{{\mathcal{D}}_{u}|u\in{\mathcal{U}}_{\mathcal{S}}\}$and target datasets ${\mathcal{D}}_{\mathcal{T}}=\{{\mathcal{D}}_{u}|u\in{\mathcal{U}}_{\mathcal{T}}\}$, respectively. Each client is trained for $E$ epochs every round. The system objectives are two-fold.
@@
where $\zeta^{*}$ is the optimal combined risk on ${\mathcal{D}}_{\mathcal{S}}\ \textrm{and}\ {\mathcal{D}}_{\mathcal{T}}$. One crucial challenge in FDG is the existence of a domain shift between the sets of source clients’ data $\mathcal{D_{\mathcal{S}}}$ and target clients $\mathcal{D_{\mathcal{T}}}$, respectively.
@@
where $\lambda$ is the scaling coefficient of domain-invariant regularization. The principle of domain-invariant learning is to learn a model that remains invariant across source domains, rather than reducing the divergence between the source and unavailable target domains. As a consequence, domain-invariant learning efficiently reduces the generalization gap between source and target domains Li et al. (2018). However, due to the decentralization of the data, each client does not have access to other clients’ data, i.e., ${\mathcal{D}}_{u}\cap{\mathcal{D}}_{v}=\varnothing,\forall u,v\in{\mathcal{U}}_{\mathcal{S}}$, $u\neq v$. Consequently, it is challenging to find the domain-invariant features within the FDG framework.
@@ 3.2 Motivation of Inter-domain Gradient Matching
Definition 1 (Invariant Gradient Direction Shi et al. (2022))
@@
Here, we abuse $g^{(r)}_{u}=g(\phi^{(r)}_{u};{\mathcal{D}}_{u})$ as the local gradient of user $u$ on round $r$ after $E$ epochs. $g(\theta;\cdot)=\theta-\theta^{(r-1)}_{g}$ is the gradient computed by learnable parameter $\theta$, thus, we do not consider the dataset. Note that Eq. (6) has a high similarity with Shi et al. (2022, Eq. (4)). Apart from (Shi et al., 2022), we use $\sum_{u\in{\mathcal{U}}_{\mathcal{S}}}\Big{\langle}g(\theta;\cdot),g(\phi^{(r)}_{u};\cdot)\Big{\rangle}$ instead of $\sum_{u\in{\mathcal{U}}_{\mathcal{S}}}\sum_{v\in{\mathcal{U}}_{\mathcal{S}}}\Big{\langle}g^{(r)}_{u},g^{(r)}_{v}\Big{\rangle}$ because $\langle g^{(r)}_{u},g^{(r)}_{v}\rangle$ is double-edged bounded by $\langle g(\theta;\cdot),g^{(r)}_{v}\rangle$ and $\langle g(\theta;\cdot),g^{(r)}_{u}\rangle$, $\forall u,v$, as explained in the following.
@@ 4.2 Inter-client Gradient Matching
$g^{(r)}_{\textrm{IGD}}=g^{(r)}_{\textrm{FL}}+\frac{\kappa\|g^{(r)}_{\textrm{FL}}\|}{\|\Gamma^{*}{\mathbf{g}}^{(r)}\|}\Gamma^{*}{\mathbf{g}}^{(r)}\quad\textrm{s.t.}\quad\Gamma^{*}=\arg\min_{\Gamma}\Gamma{\mathbf{g}}^{(r)}\cdot g^{(r)}_{\textrm{FL}}+\kappa\|g^{(r)}_{\textrm{FL}}\|\|g^{(r)}_{\Gamma}\|,$ (13)
@@ 5 Theoretical Analysis
where $d_{{\mathcal{G}}\circ\theta}(A,B)$ is the gradient divergence of model $\theta$ when training in two domains $A,B$.
@@
where $d_{{\mathcal{G}}\circ\theta}(\hat{{\mathcal{D}}}_{u},\hat{{\mathcal{D}}}_{v})$ denotes the gradient divergence of model $\theta$ when training on two different domains $\hat{{\mathcal{D}}}_{u},\hat{{\mathcal{D}}}_{v}$.
@@ 6.3 Ablation Studies
Evaluation on gradient invariance: To assess the invariant property of FedOMG, we calculate the cosine similarity between the gradients of each source domain and the global gradient, and visualize in Figs. 5a, 5b. As it can be seen from the figures, the significance of FedOMG compared to FedAvg is two-fold. First of all, the cosine similarities of all domains in FedOMG hold a low variance among domains. This mean that all of the domains share the same divergence compared with the global gradients. Secondly, the overwhelmingly high cosine value in domain C of FedAvg may results from the overfitting to domain C. As a consequence, the FedAvg significantly lacks generalization capability compared to that of the FedOMG (see Fig. 5c).
@@ 7 Conclusion
In this paper, we present FedOMG, a novel and effective federated learning (FL) domain generalization method aimed at addressing the challenge of heterogeneous client data distributions by learning an invariant gradient. Our method introduces two key innovations: (1) utilizing the local gradient as a critical factor for server-side optimization, and (2) devising a training strategy that identifies optimal coefficients to approximate the invariant gradient across client source domains. We provide theoretical proof to ensure the global invariant gradient solution and conduct extensive experiments demonstrating FedOMG’s significant improvements in both standard FL and domain generalization (FDG) settings. This work establishes an optimal approach for server optimization, leading to enhanced FL performance while maintaining privacy constraints.
@@ Appendix A Related Works
Federated Learning: To contextualize our paper’s contribution to on-server optimization, FL can be divided into two primary categories: 1) local training and 2) on-server aggregation. In the first category, substantial efforts have been dedicated to improving the performance of FL, particularly in scenarios involving non-IID data. FedRod (Chen & Chao, 2022) proposes using balanced softmax for training a generic model while utilizing vanilla softmax for personalized heads. FedBABU (Oh et al., 2022) keeps the global classifier fixed during feature representation learning and performing local adaptation through fine-tuning. FedPAC (Xu et al., 2023) adds a regularizer on local training to learn task-invariant representations. Similar to FedPAC, FedAS (Yang et al., 2024) finds the shared parameters which are task-invariant by leveraging the parameter alignment technique. Recently, researchers have focused on on-server aggregation to improve the performance of FL. This approach can be implemented by utilizing generative AI on the server with pseudo data Zhu et al. (2021), or leveraging local gradients to adjust the learning rate (Jhunjhunwala et al., 2023; Panchal et al., 2023).
@@ D.4 Model
D.5 Hyperparameters
In this section, we present the hyperparameters chosen for FedOMG across different datasets.
Table 4: Hyperparameter Summary on PACS, VLCS, and OfficeHome
Method Hyperparameter PACS VLCS OfficeHome FedOMG Global lr $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ Training lr 25 25 25 Iteration 21 21 21 Momentum 0.5 0.5 0.5 Searching radius 0.5 0.5 0.5 FedSR+OMG Global lr $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ Training lr 25 25 25 Iteration 21 21 21 Momentum 0.5 0.5 0.5 Searching radius 0.5 0.5 0.5 L2 Regularizer 0.01 0.01 0.01 Cmi Regularizer 0.001 0.001 0.001 FedIIR+OMG Hyperparameter $\gamma$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ Global lr $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ Training lr 25 25 25 Iteration 21 21 21 Momentum 0.5 0.5 0.5 Searching radius 0.5 0.5 0.5 FedSAM+OMG Global lr $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ Training lr 25 25 25 Iteration 21 21 21 Momentum 0.5 0.5 0.5 Searching radius 0.5 0.5 0.5 Perturbation control $\rho$ 0.6 0.6 0.6 momentum parameter $\beta$ 0.6 0.6 0.6
@@ D.6 Implementation Details
To evaluate the performance of our proposed federated learning (FL) method, we adopt the methodologies described in Lin et al. (2020); Acar et al. (2021) to simulate non-IID data. We utilize the Dirichlet distribution with two levels of data heterogeneity, specifically $\alpha=0.1$ and $\alpha=0.5$. Our experimental setup comprises three configurations of 100 clients, with join rate ratios of 1, 0.6, and 0.4, respectively, and global communication rounds set at 800. Each client undergoes 5 local training rounds using a local learning rate of 0.005, an SGD optimizer, and a batch size of 16. To ensure a fair comparison, all methods are evaluated with the same network architecture and settings.
For FDG scenarios, we evaluate model performance using the conventional leave-one-domain-out method Guo et al. (2023); Fan et al. (2024), where one domain is designated as the test domain and all other domains are used for training. Number of clients is set to match the number of source domains. In line with Zhang et al. (2023b); Park et al. (2023); Guo et al. (2023), we use a pretrained ResNet-18 backbone model. Our experiments involve 100 global communication rounds with 5 local training rounds, utilizing the SGD optimizer with a learning rate of 0.001 and a training batch size of 16. Each experiment is conducted three times, and we report the average performance of the global model on the test domain, with each domain serving as the test domain once.

### v2

# Federated Domain Generalization with Data-free On-server Matching Gradient

@@ Abstract
Domain Generalization (DG) aims to learn from multiple known source domains a model that can generalize well to unknown target domains. One of the key approaches in DG is training an encoder which generates domain-invariant representations. However, this approach is not applicable in Federated Domain Generalization (FDG), where data from various domains are distributed across different clients. In this paper, we introduce a novel approach, dubbed Federated Learning via On-server Matching Gradient (FedOMG), which can efficiently leverage domain information from distributed domains. Specifically, we utilize the local gradients as information about the distributed models to find an invariant gradient direction across all domains through gradient inner product maximization. The advantages are two-fold: 1) FedOMG can aggregate the characteristics of distributed models on the centralized server without incurring any additional communication cost, and 2) FedOMG is orthogonal to many existing FL/FDG methods, allowing for additional performance improvements by being seamlessly integrated with them. Extensive experimental evaluations on various settings demonstrate the robustness of FedOMG compared to other FL/FDG baselines. Our method outperforms recent SOTA baselines on four FL benchmark datasets (MNIST, EMNIST, CIFAR-10, and CIFAR-100), and three FDG benchmark datasets (PACS, VLCS, and OfficeHome). The reproducible code is publicly available  [^1] https://github.com/skydvn/fedomg.
@@ 1 Introduction
Federated Learning (FL) has gained widespread recognition due to its ability to train models collaboratively across multiple clients while keeping their individual data secure. However, one practical challenge, how to ensure that models trained on sites with heterogeneous distributions generalize to target clients with unknown distributions, known as Federated Domain Generalization (FDG), remains under-explored. While label distribution shift has been considered in traditional FL, FDG focuses on the feature shift among clients and considers each client as an individual domain.
FDG shares a similar goal as standard Domain Generalization (DG) (Nguyen et al., 2022b; Li et al., 2018), i.e., generalizing from multi-source domains to unseen domains. However, unlike DG, where knowledge from various domains can be jointly utilized to develop an efficient algorithm, FDG prohibits direct data sharing among clients, which makes most existing DG methods hardly applicable. To address the challenges inherent to FDG, recent methods circumvent these difficulties by adopting alternative strategies, e.g., unbiased local training within each isolated domain, local data preprocessing (Huang et al., 2022; 2024), or knowledge sharing among users (Liu et al., 2021; Chen et al., 2023). As a consequence, an open question remains for FDG:
@@
    Due to the design of on-server gradient matching, our method is orthogonal to many classic FDG methods, where current FDG methods focus on the local training (Huang et al., 2022; 2024). As a consequence, the integration of FedOMG with other FDG methods can achieve improved performance.
@@ 2 Preliminaries
Problem Setup: We consider an FL system comprising a set of source clients ${\mathcal{U}}_{\mathcal{S}}=\{u|u=1,2,\ldots,U_{\mathcal{S}}\}$, and target clients ${\mathcal{U}}_{\mathcal{T}}=\{u|u=U_{\mathcal{S}}+1,U_{\mathcal{S}}+2,\ldots,U_{\mathcal{S}}+U_{\mathcal{T}}\}$. Each client $u$ gains access to its data ${\mathcal{D}}_{u}=\{x_{i},y_{i}\}^{N_{u}}_{i=1}$ with $N_{u}$ data samples, where $u\in\{{\mathcal{U}}_{\mathcal{S}};{\mathcal{U}}_{\mathcal{T}}\}$. Source and target clients are assigned to the source datasets ${\mathcal{D}}_{\mathcal{S}}=\{{\mathcal{D}}_{u}|u\in{\mathcal{U}}_{\mathcal{S}}\}$ and target datasets ${\mathcal{D}}_{\mathcal{T}}=\{{\mathcal{D}}_{u}|u\in{\mathcal{U}}_{\mathcal{T}}\}$, respectively. Each client is trained for $E$ epochs every round. The system objectives are two-fold.
@@
where $\zeta^{*}$ is the optimal combined risk on ${\mathcal{D}}_{\mathcal{S}}\ \textrm{and}\ {\mathcal{D}}_{\mathcal{T}}$. One crucial challenge in FDG is the existence of a domain shift between the sets of source clients’ data $\mathcal{D_{\mathcal{S}}}$ and target clients’ data $\mathcal{D_{\mathcal{T}}}$, respectively.
@@
where $\lambda$ is the scaling coefficient of domain-invariant regularization. The principle of domain-invariant learning is to learn a model that remains invariant across source domains, rather than reducing the divergence between the source and unavailable target domains. As a consequence, domain-invariant learning efficiently reduces the generalization gap between source and target domains (Li et al., 2018). However, due to the decentralization of the data, each client does not have access to other clients’ data, i.e., ${\mathcal{D}}_{u}\cap{\mathcal{D}}_{v}=\varnothing,\forall u,v\in{\mathcal{U}}_{\mathcal{S}}$, $u\neq v$. Consequently, it is challenging to find the domain-invariant features within the FDG framework.
@@ 3.2 Motivation of Inter-domain Gradient Matching
Definition 1 (Invariant Gradient Direction (Shi et al., 2022))
@@
Here, we abuse $g^{(r)}_{u}=g(\phi^{(r)}_{u};{\mathcal{D}}_{u})$ as the local gradient of user $u$ on round $r$ after $E$ epochs. $g(\theta;\cdot)=\theta-\theta^{(r-1)}_{g}$ is the gradient computed by learnable parameter $\theta$, thus, we do not consider the dataset. Note that Eq. (6) has a high similarity with (Shi et al., 2022, Eq. (4)). Apart from (Shi et al., 2022), we use $\sum_{u\in{\mathcal{U}}_{\mathcal{S}}}\Big{\langle}g(\theta;\cdot),g(\phi^{(r)}_{u};\cdot)\Big{\rangle}$ instead of $\sum_{u\in{\mathcal{U}}_{\mathcal{S}}}\sum_{v\in{\mathcal{U}}_{\mathcal{S}}}\Big{\langle}g^{(r)}_{u},g^{(r)}_{v}\Big{\rangle}$ because $\langle g^{(r)}_{u},g^{(r)}_{v}\rangle$ is double-edged bounded by $\langle g(\theta;\cdot),g^{(r)}_{v}\rangle$ and $\langle g(\theta;\cdot),g^{(r)}_{u}\rangle$, $\forall u,v$, as explained in the following.
@@ 4.2 Inter-client Gradient Matching
$g^{(r)}_{\textrm{IGD}}=g^{(r)}_{\textrm{FL}}+\frac{\kappa\|g^{(r)}_{\textrm{FL}}\|}{\|\Gamma^{*}{\mathbf{g}}^{(r)}\|}\Gamma^{*}{\mathbf{g}}^{(r)}\quad\textrm{s.t.}\quad\Gamma^{*}=\arg\min_{\Gamma}\Gamma{\mathbf{g}}^{(r)}\cdot g^{(r)}_{\textrm{FL}}+\kappa\|g^{(r)}_{\textrm{FL}}\|\|g^{(r)}_{\Gamma}\|.$ (13)
@@ 5 Theoretical Analysis
where $d_{{\mathcal{G}}\circ\theta}(A,B)$ is the gradient divergence of model $\theta$ when training in two domains $A$ and $B$.
@@
where $\hat{{\mathcal{D}}}_{u},\hat{{\mathcal{D}}}_{v}$ are the sampled counterparts from the domain ${\mathcal{D}}_{u},{\mathcal{D}}_{u}$, respectively. $d_{{\mathcal{G}}\circ\theta}(\hat{{\mathcal{D}}}_{u},\hat{{\mathcal{D}}}_{v})$ denotes the gradient divergence of model $\theta$ when training on two different domains $\hat{{\mathcal{D}}}_{u}$ and $\hat{{\mathcal{D}}}_{v}$.
@@ 6.3 Ablation Studies
Evaluation on gradient invariance: To assess the invariant property of FedOMG, we calculate the cosine similarity between the gradients of each source domain and the global gradient, and visualize in Figs. 5a, 5b. As it can be seen from the figures, the significance of FedOMG compared to FedAvg is two-fold. First of all, the cosine similarities of all domains in FedOMG hold a low variance among domains. This mean that all of the domains share the same divergence compared with the global gradients. Secondly, the overwhelmingly high cosine value in domain S of FedAvg may results from the overfitting to domain S. As a consequence, the FedAvg significantly lacks generalization capability compared to that of the FedOMG (see Fig. 5c).
@@ 7 Conclusion
In this paper, we present FedOMG, a novel and effective federated domain generalization method aimed at addressing the challenge of heterogeneous client data distributions by learning an invariant gradient. Our method introduces two key innovations: (1) utilizing the local gradient as a critical factor for server-side optimization, and (2) devising a training strategy that identifies optimal coefficients to approximate the invariant gradient across client source domains. We provide theoretical proof to ensure the global invariant gradient solution and conduct extensive experiments demonstrating FedOMG’s significant improvements in both standard FL and domain generalization (FDG) settings. This work establishes an optimal approach for server optimization, leading to enhanced FL performance while maintaining privacy constraints.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) under the Artificial Intelligence Convergence Innovation Human Resources Development (IITP-2025-RS-2023-00254177) grant funded by the Korea government(MSIT). This research was supported by the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program(IITP-2025-RS-2023-00260098) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation). This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00336962). The work of Quoc-Viet Pham is supported in part by the European Union under the ENSURE-6G project (Grant Agreement No. 101182933).
@@ Appendix A Related Works
Federated Learning: To contextualize our paper’s contribution to on-server optimization, FL can be divided into two primary categories: 1) local training and 2) on-server aggregation. In the first category, substantial efforts have been dedicated to improving the performance of FL, particularly in scenarios involving non-IID data. FedRod (Chen & Chao, 2022) proposes using balanced softmax for training a generic model while utilizing vanilla softmax for personalized heads. FedBABU (Oh et al., 2022) keeps the global classifier fixed during feature representation learning and performing local adaptation through fine-tuning. FedPAC (Xu et al., 2023) adds a regularizer on local training to learn task-invariant representations. Similar to FedPAC, FedAS (Yang et al., 2024) finds the shared parameters which are task-invariant by leveraging the parameter alignment technique. Recently, researchers have focused on on-server aggregation to improve the performance of FL. This approach can be implemented by utilizing generative AI on the server with pseudo data (Zhu et al., 2021), or leveraging local gradients to adjust the learning rate (Jhunjhunwala et al., 2023; Panchal et al., 2023).
@@ D.4 Model
D.5 Hyper-parameters
In this section, we present the hyper-parameters chosen for FedOMG across different datasets.
Table 4: Hyper-parameter Summary on PACS, VLCS, and OfficeHome
Method Hyper-parameter PACS VLCS OfficeHome FedOMG Global lr $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ Training lr 25 25 25 Iteration 21 21 21 Momentum 0.5 0.5 0.5 Searching radius 0.5 0.5 0.5 FedSR+OMG Global lr $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ Training lr 25 25 25 Iteration 21 21 21 Momentum 0.5 0.5 0.5 Searching radius 0.5 0.5 0.5 L2 Regularizer 0.01 0.01 0.01 Cmi Regularizer 0.001 0.001 0.001 FedIIR+OMG Hyper-parameter $\gamma$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ Global lr $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ Training lr 25 25 25 Iteration 21 21 21 Momentum 0.5 0.5 0.5 Searching radius 0.5 0.5 0.5 FedSAM+OMG Global lr $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ Training lr 25 25 25 Iteration 21 21 21 Momentum 0.5 0.5 0.5 Searching radius 0.5 0.5 0.5 Perturbation control $\rho$ 0.6 0.6 0.6 momentum parameter $\beta$ 0.6 0.6 0.6
@@ D.6 Implementation Details
To evaluate the performance of our proposed FDG method, we adopt the methodologies described in Lin et al. (2020); Acar et al. (2021) to simulate non-IID data. We utilize the Dirichlet distribution with two levels of data heterogeneity, specifically $\alpha=0.1$ and $\alpha=0.5$. Our experimental setup comprises three configurations of 100 clients, with join rate ratios of 1, 0.6, and 0.4, respectively, and global communication rounds set at 800. Each client undergoes 5 local training rounds using a local learning rate of 0.005, an SGD optimizer, and a batch size of 16. To ensure a fair comparison, all methods are evaluated with the same network architecture and settings.
For FDG scenarios, we evaluate model performance using the conventional leave-one-domain-out method (Guo et al., 2023; Fan et al., 2024), where one domain is designated as the test domain and all other domains are used for training. Number of clients is set to match the number of source domains. In line with Zhang et al. (2023b); Park et al. (2023); Guo et al. (2023), we use a pretrained ResNet-18 backbone model. Our experiments involve 100 global communication rounds with 5 local training rounds, utilizing the SGD optimizer with a learning rate of 0.001 and a training batch size of 16. Each experiment is conducted three times, and we report the average performance of the global model on the test domain, with each domain serving as the test domain once.